# CSI-BERT2: A BERT-inspired Framework for Efficient CSI Prediction and Classification in Wireless Communication and Sensing

### v2

@@ IV-C3 CSI Classification Task
The robust and powerful capabilities of our CSI-BERT2 model in CSI classification tasks are demonstrated in Table III, where it achieves the best performance across all scenarios. Notably, the original CSI-BERT fails to outperform models such as Chen et al. [51] and WiGrumt [49] in many scenarios. However, our CSI-BERT2 shows substantial improvement due to the enhanced model structure, thanks to the ARL capturing the relationships and importance among subcarriers, and the MLP-based time embedding layer that encodes time information more effectively. Moreover, aided by the time embedding mechanism, CSI-BERT2 can be effectively applied in scenarios where the training data includes samples with varying sampling rates, or where the sampling rate of the testing set differs from that of the training set. This is a practical feature, especially in federated learning scenarios, where we cannot guarantee that all clients have the same sampling rate. In Table V, we compare the performance of various models under three conditions: (1) both training and testing sets contain 100 Hz CSI, (2) both contain 50 Hz CSI, and (3) the training set contains only 100 Hz CSI while the testing set contains 50 Hz CSI. For fairness, we provide the CSI processed by linear interpolation for the other non-CSI-BERT models. It can be seen that other models are significantly influenced by the different sampling rates, with some even failing to function properly in particular scenarios. In contrast, CSI-BERT2 is almost unaffected by any impact, thanks to the help of time embedding.
@@
TABLE V: CSI Classification Performance under Different Sampling Rate
Data Training Set: 100Hz+50Hz; Testing Set: 100Hz+50Hz Model MLP CNN RNN [50] LSTM [48] Chen et al. [51] WiGRUNT [49] CSI-BERT [7] CSI-BERT2 Gesture Recognition 16.51% 14.54% 15.13% 17.37% 16.59% 74.47% 64.61% 97.04% People Identification 13.47% 15.52% 13.41% 13.47% 13.72% 81.25% 70.83% 99.54% Action Recognition 70.97% 67.67% 70.02% 59.63% 75.18% 66.78% 78.18% 88.35% Fall Detection 81.10% 75.80% 83.92% 84.83% 85.98% 80.72% 92.98% 93.64% People Number Estimation 84.03% 46.82% 81.25% 82.00% 80.09% 76.13% 86.93% 92.54% Data Training Set: 100Hz; Testing Set: 50Hz Model MLP CNN RNN [50] LSTM [48] Chen et al. [51] WiGRUNT [49] CSI-BERT [7] CSI-BERT2 Gesture Recognition 69.79% 20.38% 36.25% 27.15% 74.89% 71.37% 79.96% 97.81% People Identification 87.29% 11.85% 82.29% 22.32% 87.22% 85.24% 94.44% 99.38% Action Recognition 68.97% 51.96% 68.07% 60.26% 76.56% 73.21% 84.56% 88.53% Fall Detection 80.83% 76.13% 80.17% 84.38% 78.61% 77.08% 94.19% 94.32% People Number Estimation 82.22% 77.03% 84.43% 42.22% 68.89% 71.11% 89.10% 94.77%
@@ IV-D1 Effect of Unsupervised Training
To illustrate the benefits of our unsupervised training mechanism for downstream tasks, we compare the model’s performance with and without unsupervised training, as shown in Fig. 6 and Table VI. While unsupervised training contributes to performance gains in the CSI classification task, the improvement remains relatively modest. However, in the CSI prediction task, unsupervised training proves to be highly beneficial, this is because the prediction task is more closely related to the mask recovery task.
@@
TABLE VI: CSI-BERT2 Performance in CSI Prediction Task
Dataset w/ Unsupervised Training w/o Unsupervised Training
@@
CommPre 1.5689 ×$\times 10^{-8}$ 0.2959 0.2319 2.0080 ×$\times 10^{-8}$ 0.3198 0.2477
@@
IV-D2 Effect of Changes on Original BERT Model
@@
Next, we compare the attention patterns of the proposed CSI-BERT models with the original BERT. As shown in Fig. 8, we illustrate the mean of attention patterns across all attention heads. The attention patterns of CSI-BERT and CSI-BERT2 consist of multiple diagonals from left to right, which is interpretable. This is because in a short time interval, CSI always changes in a particular regular pattern. Additionally, it is reasonable that their attention patterns are approximately symmetric, which corresponds to the mechanism of bi-directional attention in BERT. However, we cannot find any regular patterns in the attention patterns of the original BERT.
@@
Furthermore, we use a gradient-weighted class activation mapping (Grad-CAM) [59] method to provide explanations for the CSI recovery task of CSI-BERT and BERT. We manually delete a given packet and use the models to recover them. After that, we calculate the Grad-CAM based on the gradient of the original input. We then combine it with the original input, as illustrated in Fig. 9, to visualize the regions the model primarily attends to during the recovery of the missing packet. It can be seen that our CSI-BERT and CSI-BERT2 models mainly focus on packets nearby the lost packet, which is consistent with our intuition. However, the original BERT only focuses on the lost packet, which is actually meaningless because the lost packet is replaced by [MASK] and can hardly provide useful information, except that the packet at this position is lost.
@@
Finally, by comparing the attention maps and Grad-CAM between CSI-BERT and CSI-BERT2, we can explain why CSI-BERT2 exhibits higher generalization. In Fig. 9, we observe that CSI-BERT2 has a more centralized activation map, suggesting it utilizes less information for training while achieving better performance than CSI-BERT. This allows CSI-BERT2 to learn a simpler attention map, as shown in Fig. 8, enabling it to focus on more robust features rather than overfitting to specific relationships in the training set. This improved generalization may be attributed to the ARL and time embedding layers, which help capture more information in the dimensions of subcarrier and time (packet).
@@
(a) BERT [18] (202020-th packet)
(b) CSI-BERT [7] (202020-th packet)
(c) CSI-BERT2 (202020-th packet)
(d) BERT [18] (808080-th packet)
(e) CSI-BERT [7] (808080-th packet)
(f) CSI-BERT2 (808080-th packet)
Figure 9: Grad-CAM: The two rows of heatmaps illustrate the Grad-CAM when we manually delete a packet at position 20 and position 80. The blank positions represent the actual lost CSI during transmission. We calculate Grad-CAM solely based on the loss function at the positions where we manually delete the packets.
@@ V Conclusion
The CSI-BERT2 model can be further applied in other CSI-related scenarios and time-series tasks. Moreover, we aim to explore how to leverage multi-source heterogeneous CSI data to train our CSI-BERT2 simultaneously. By doing so, we hope to replicate the generalization capabilities of large language models (LLMs) and large vision models (LVMs) in the context of wireless sensing.

### v3

@@ IV-C3 CSI Classification Task
The robust and powerful capabilities of our CSI-BERT2 model in CSI classification tasks are demonstrated in Table V, where it achieves the best performance across all scenarios. Notably, the original CSI-BERT fails to outperform models such as Chen et al. [56] and WiGrumt [54] in many scenarios. However, our CSI-BERT2 shows substantial improvement due to the enhanced model structure, thanks to the ARL capturing the relationships and importance among subcarriers, and the MLP-based time embedding layer that encodes time information more effectively. Moreover, aided by the time embedding mechanism, CSI-BERT2 can be effectively applied in scenarios where the training data includes samples with varying sampling rates, or where the sampling rate of the testing set differs from that of the training set. This is a practical feature, especially in federated learning scenarios, where we cannot guarantee that all clients have the same sampling rate. In Table VII, we compare the performance of different models where the training and testing sets both have 100Hz and 50Hz CSI, as well as where the training set only has 100Hz CSI but the testing set has 50Hz CSI. For fairness, we provide the CSI processed by linear interpolation for the other non-CSI-BERT models. It can be seen that other models are significantly influenced by the different sampling rates, with some even failing to function properly in particular scenarios. In contrast, CSI-BERT2 is almost unaffected by any impact, thanks to the help of time embedding.
@@
TABLE VII: CSI Classification Performance under Different Sampling Rate
Data Training Set: 100Hz+50Hz; Testing Set: 100Hz+50Hz Model MLP CNN RNN [55] LSTM [53] Chen et al. [56] WiGRUNT [54] CSI-BERT [11] CSI-BERT2 Gesture Recognition 16.51% 14.54% 15.13% 17.37% 16.59% 74.47% 64.61% 97.04% Person Identification 13.47% 15.52% 13.41% 13.47% 13.72% 81.25% 70.83% 99.54% Action Recognition 70.97% 67.67% 70.02% 59.63% 75.18% 66.78% 78.18% 88.35% Fall Detection 81.10% 75.80% 83.92% 84.83% 85.98% 80.72% 92.98% 93.64% People Number Estimation 84.03% 46.82% 81.25% 82.00% 80.09% 76.13% 86.93% 92.54% Data Training Set: 100Hz; Testing Set: 50Hz Model MLP CNN RNN [55] LSTM [53] Chen et al. [56] WiGRUNT [54] CSI-BERT [11] CSI-BERT2 Gesture Recognition 69.79% 20.38% 36.25% 27.15% 74.89% 71.37% 79.96% 97.81% Person Identification 87.29% 11.85% 82.29% 22.32% 87.22% 85.24% 94.44% 99.38% Action Recognition 68.97% 51.96% 68.07% 60.26% 76.56% 73.21% 84.56% 88.53% Fall Detection 80.83% 76.13% 80.17% 84.38% 78.61% 77.08% 94.19% 94.32% People Number Estimation 82.22% 77.03% 84.43% 42.22% 68.89% 71.11% 89.10% 94.77%
@@ IV-D1 Effect of Unsupervised Training
To illustrate the benefits of our unsupervised training mechanism for downstream tasks, we compare the model’s performance with and without unsupervised training, as shown in Fig. 6 and Table VIII. While unsupervised training contributes to performance gains in the CSI classification task, the improvement remains relatively modest. However, in the CSI prediction task, unsupervised training proves to be highly beneficial, this is because the prediction task is more closely related to the mask recovery task.
@@
TABLE VIII: CSI-BERT2 Performance in CSI Prediction Task
Dataset Proposed w/o Unsupervised Training
@@
CommPre 1.5689 $\times 10^{-8}$ 0.2959 0.2319 2.0080 $\times 10^{-8}$ 0.3198 0.2477
@@
Dataset w/ Time Embedding w/o ARL
MSE SMAPE MAPE MSE SMAPE MAPE
CommPre $1.7676\times 10^{-8}$ 0.3054 0.2329 $1.6995\times 10^{-8}$ 0.3062 0.2123
WiGesture [11] 3.8399 0.1645 0.1450 5.5027 0.2075 0.1678
WiFall [41] 5.4054 0.1573 0.1434 5.2610 0.1515 0.1382
WiCount 5.9274 0.1843 0.1550 5.7969 0.1829 0.1558
IV-D2 Effect of Changes on BERT and CSI-BERT
@@
We then evaluate the efficacy of the introduced time embedding layer and the ARL for capturing temporal and spatial information, respectively, in comparison to CSI-BERT. The results of the ablation study for CSI classification and CSI prediction are shown in Fig. 6 and Table VIII, respectively. For the CSI classification task, we observed that the time embedding layer and ARL led to significant improvements in the WiFall and WiCount datasets, but not in the WiGesture dataset. This discrepancy may be related to the difficulty of the tasks: in the WiGesture dataset, the proposed method achieved nearly 100% accuracy even without pre-training, ARL, or the time embedding layer. In contrast, the proposed modules demonstrate greater benefits for the more complex tasks in the WiFall and WiCount datasets. For the CSI prediction task, both ARL and the time embedding layer significantly reduced prediction errors across all datasets. These findings illustrate that the two modules effectively help the model capture the temporal and spatial relationships within the data.
Next, we compare the attention patterns of the CSI-BERT models with those of the original BERT. Specifically, we calculate the mean value of the attention matrix from each layer:
$\overline{A}=\frac{1}{LH}\sum_{l=1}^{L}\sum_{h=1}^{H}\text{Softmax}\left(\frac{{Q_{l,h}}{K_{l,h}}^{T}}{\sqrt{d_{k}}}\right)\ ,$ (14)
@@
where $Q_{l,h}$} and $K_{l,h}$} represent the query and key of the $h$-th head in the $l$-th layer, $L$ and $H$ denote the number of layers and heads, respectively, and $d_{k}$ indicates their feature dimension. As shown in Fig. 8, we illustrate the mean of attention patterns across all attention heads. The attention patterns of CSI-BERT and CSI-BERT2 consist of multiple diagonals from left to right, which is interpretable. This is because in a short time interval, CSI always changes in a particular regular pattern. Additionally, it is reasonable that their attention patterns are approximately symmetric, which corresponds to the mechanism of bi-directional attention in BERT. However, we cannot find any regular patterns in the attention patterns of the original BERT.
Furthermore, we utilize a Saliency Map to provide explanations for the CSI recovery task of CSI-BERT and BERT. We manually delete a given packet and use the models to recover it. Subsequently, we calculate the Saliency Map based on the gradient of the original input, as given by $|\frac{\partial(y-\hat{y})^{2}}{\partial x}|$, where $x$, $y$, and $\hat{y}$ represent the input, ground truth, and output, respectively.
@@
We then combine it with the original input, as illustrated in Fig. 9, to visualize the regions the model primarily attends to during the recovery of the missing packet. It can be seen that our CSI-BERT and CSI-BERT2 models mainly focus on packets nearby the lost packet, which is consistent with our intuition. However, the original BERT only focuses on the lost packet, which is actually meaningless because the lost packet is replaced by [MASK] and can hardly provide useful information, except that the packet at this position is lost.
@@
Finally, by comparing the attention maps and Saliency Map between CSI-BERT and CSI-BERT2, we can explain why CSI-BERT2 exhibits higher generalization. In Fig. 9, we observe that CSI-BERT2 has a more centralized activation map, suggesting it utilizes less information for training while achieving better performance than CSI-BERT. This allows CSI-BERT2 to learn a simpler attention map, as shown in Fig. 8, enabling it to focus on more robust features rather than overfitting to specific relationships in the training set. This improved generalization may be attributed to the ARL and time embedding layers, which help capture more information in the dimensions of subcarrier and time (packet).
@@
(a) BERT[22] (2020-th packet)
(b) CSI-BERT[11] (2020-th packet)
(c) CSI-BERT2 (2020-th packet)
(d) BERT [22] (8080-th packet)
(e) CSI-BERT [11] (8080-th packet)
(f) CSI-BERT2 (8080-th packet)
Figure 9: Saliency Map: The two rows of heatmaps illustrate the Saliency Map when we manually delete a packet at position 20 and position 80. The blank positions represent the actual lost CSI during transmission. We calculate Saliency Map solely based on the loss function at the positions where we manually delete the packets.
@@ V Conclusion
Additionally, several aspects of our method warrant further exploration in future work. First, this paper’s experiments focus solely on a single-antenna and single pair TX-RX scenario. Future research could expand CSI-BERT2 to more general scenarios and explore better encoding methods to process multiple antenna inputs. Second, data heterogeneity poses a significant challenge in CSI sensing and prediction. It would be worthwhile to investigate whether CSI-BERT2 can be modified to address this challenge and serve as a foundation model, given its demonstrated potential across multiple tasks and benefits from pre-training.
Biography
[Uncaptioned image] Zijian Zhao received the B.Eng. degree in computer science and technology from School of Computer Since and Engineering, Sun Yat-sen University in 2024. He is currently pursuing the Ph.D. degree in civil engineering (scientific computation) with Department of Civil and Environmental Engineering, The Hong Kong University of Science and Technology. He was a visiting student in Shenzhen Research Institute of Big Data, The Chinese University of Hong Kong (Shenzhen) from 2023 to 2024. His current research interests include deep learning, reinforcement learning, intelligent transportation, mobile computing, and wireless sensing.
[Uncaptioned image] Fanyi Meng received the B.Eng. degree in communication engineering from Southern University of Science and Technology in 2023. He is currently pursuing the M.Phil degree with Computer and Information Engineering, The Chinese University of Hong Kong (Shenzhen). His current research interests include Integrated Sensing and Communication, wireless simulation and digital twin.
[Uncaptioned image] Zhonghao Lyu received the Ph.D. degree in computer and information engineering from The Chinese University of Hong Kong, Shenzhen, China, in 2024, the M.Eng. degree in information and communication engineering from the University of Science and Technology of China, Hefei, China, in 2021, and the B.Eng. degree from Dalian University of Technology, Dalian, China, in 2018. From 2022 to 2024, he was a visiting student with the Shenzhen Research Institute of Big Data. He is currently a Post-Doctoral Research Fellow with the School of Electrical Engineering and Computer Science, KTH Royal Institute of Technology, Sweden. His research interests include large AI models, edge intelligence, and UAV communications.
[Uncaptioned image] Hang Li received the B.E. and M.S. degrees from Beihang University, Beijing, China, in 2008 and 2011, respectively, and the Ph.D. degree from Texas A&M University, College Station, TX, USA, in 2016. He was a postdoctoral research associate with Texas A&M and University of California-Davis (Sept. 2016-Mar. 2018). After being a visiting research scholar (Apr. 2018 – June 2019) at Shenzhen Research Institute of Big Data, Shenzhen, China, he has been a research scientist since June 2019. His current research interests include wireless networks, Internet of things, stochastic optimization, and applications of machine learning. He is recognized as Overseas High-Caliber Personnel (Level C) at Shenzhen in 2020.
[Uncaptioned image] Xiaoyang Li (Member, IEEE) received the B.Eng. degree from the Southern University of Science and Technology (SUSTech) in 2016 and the Ph.D. degree from The University of Hong Kong in 2020. From 2020 to 2022, he was a Presidential Distinguished Research Fellow at SUSTech. He is currently an Associate Research Fellow with the Shenzhen Research Institute of Big Data, and an Adjunct Assistant Professor with the Chinese University of Hong Kong, Shenzhen. He is a recipient of the Youth Talent Support Program by China Communication Society, Forbes China 30 under 30, Overseas Youth Talent in Guangdong, Overseas High-caliber Personnel in Shenzhen, Outstanding Research Fellow in Shenzhen, the Best Paper Award of IEEE 4th International Symposium on Joint Communications & Sensing, and the Exemplary Reviewer of IEEE Wireless Communications Letters. He served as the Editor of Journal of Information and Intelligence, the Workshop/Session Chairs of IEEE ICASSP/ICC/WCNC/PIMRC/MIIS. His research interests include integrated sensing-communication-computation, edge learning, and network optimization.
[Uncaptioned image] Guangxu Zhu (Member, IEEE) received the Ph.D. degree in electrical and electronic engineering from The University of Hong Kong in 2019. Currently he is a senior research scientist and deputy director of network system optimization center at the Shenzhen research institute of big data, and an adjunct associate professor with the Chinese University of Hong Kong, Shenzhen. His recent research interests include edge intelligence, semantic communications, and integrated sensing and communication. He is a recipient of the 2023 IEEE ComSoc Asia-Pacific Best Young Researcher Award and Outstanding Paper Award, the World’s Top 2% Scientists by Stanford University, the ”AI 2000 Most Influential Scholar Award Honorable Mention”, the Young Scientist Award from UCOM 2023, the Best Paper Award from WCSP 2013 and IEEE JSnC 2024. He serves as associate editors at top-tier journals in IEEE, including IEEE TMC, TWC and WCL. He is the vice co-chair of the IEEE ComSoc Asia-Pacific Board Young Professionals Committee.